OpenStudy (anonymous):

Help.. 500 decreased by 55% = please show me how to work.

OpenStudy (anonymous):

500-55%of of 500=500-55/100 X 500=225

OpenStudy (anonymous):

Hey , 500 decreased by 55% = 500 - ( 500 * 55/100)

OpenStudy (anonymous):

first find 55% of 500 then say 500 - 55%of500 55% of 500 = 500 * 55/100 = 275 500 - 275 = 225

OpenStudy (anonymous):

or you can do do in this way too 55% decrease means 45% is left Therefore 500 decreased by 55% =45% of 500=225

OpenStudy (anonymous):

can you show me using proportion, i still do not see how to do this.

OpenStudy (anonymous):



OpenStudy (anonymous):

nw \[\frac{ 100 }{ 45 }\]=\[\frac{ 500 }{ x }\]=>x=225

OpenStudy (anonymous):

Have u got it friend @wildwoman???

OpenStudy (anonymous):

Thank you so much, I get it. I just needed to set up in a waty I can work with .

OpenStudy (anonymous):

:)